An electron has an uncertainty in its position of 603pm .

What is the uncertainty in its velocity?
Express your answer with the appropriate units.
Answer #1

\Deltap\Delta x= h/4\pi

m\Deltav\Deltax = h/4\pi

\Deltav = h/4\pi* m\Deltax

h=6.63 *10-34 m2 Kg sec-1   mass of electron (m) = 9.1*10-31 kg   \Deltax = 603pm = 603 *10-12 m

\Deltav = h/4\pi* m\Deltax

      = 6.63 *10-34 m2 Kg sec-1 / 4 * 3.14 * 9.1*10-31 kg * 603*10-12 m

      = 9.62 *104 msec-1